# Information Discovery in e-Commerce

### v1

@@ 5.5.2 Conversational e-commerce search
Originating from early studies on interactive information retrieval [croft1987i3r, belkin1995cases], conversational search refers to the process of interacting with a dialogue system to search for information. Conversational search allows users to express their information needs by directly conducting conversations with search engines. Unlike traditional query-based search engines, conversational search systems capture users’ intent by taking advantage of the flexibility of mixed-initiative interactions and by providing useful information more directly using human-like responses [10.1145/3020165.3020183, Vtyurina:2017:ECS:3027063.3053175, ren2021wizard, vakulenko2021large]. User studies have been conducted to study whether conversational search is needed and what it should look like. Trippas:2018:IDS:3176349.3176387 conducted a laboratory-based observational study, where pairs of people perform search tasks communicating verbally. The authors found that conversation search paradigms are more complex and interactive than traditional search scenarios. Moreover, it is difficult to simulate human-human interactions in a conversational search session. To address this problem, ren2021wizard collected a dataset of human-human dialogues about conversational search in a wizard-of-oz fashion, namely wizard of search engine (WISE), where two workers play the role of seeker and intermediary, respectively.
@@
In a conversational search session, the user first initializes the conversation with a request, then the conversational agent iteratively asks the user about their preferences and estimates user interest based on their feedback. At last, the agent retrieves the information and generates the response. Many studies focus on the second stage, including generating clarifying queries [Hamedwww2020, liu2021learning, ghanem2022question] and understanding users’ intent [wu-yan-2018-deep, gao2021advances, TRIPPAS2020102162]. radlinski-2017-theoretical considered the question of what properties would be desirable for a system so that the system enables users to answer a variety of information needs in a natural and efficient manner. azzopardi2018conceptualizing outlined the actions and intents of users and systems and explain how these actions enable users to explore the search space and resolve their information needs. ghanem2022question proposed a method to generate user queries for story-based reading comprehension skills. Recently, the response generation problem in conversational search has been addressed. ren2021wizard proposed a modular end-to-end neural architecture to transfer the output from the intent understanding to improve the response generation. ye2022structured designed an interconnected network to co-generate structured and natural responses that allow for bidirectional semantic associations to generate responses.
@@
In e-commerce, conversational search systems can help consumers access products through instant conversational interactions on mobile phones or other smart devices. Also, conversational e-commerce search alleviates the burden of reformulating queries and browsing through dozens of products in e-commerce search [zou2022learning]. Moreover, conversational e-commerce search provides a more natural way to collect explicit feedback from users to understand their preferences [zamani2022conversational]. As an early attempt, zhang2018towards proposed a paradigm for conversational product search to ask users about their preferred values of an aspect and adopted a memory network to retrieve search results. Based on that, bi2019conversational focused on product-seeking conversations and proposed an aspect-value likelihood model for negative feedback, with a multivariate Bernoulli distribution to generate explainable e-commerce aspects. To improve the quality of representation in conversational product search, zou2022learning integrated the representation learning of user, query, item, and conversation into a unified generative framework.
@@
There are several open questions and research directions for future work in conversational e-commerce search. First, obtaining data in real-world scenarios is still a challenging problem. Several studies, including, e.g., ConvPS [zou2022learning], have applied simulated data in the experiments. It would be useful to extend the observational experiment into a wizard-of-oz setting by establishing a real-world dataset in the future. We also observe that the ongoing trend in simulating users in task-oriented dialogues could provide more useful insights [sunzhang2021simulator]. Second, modeling user-system interactions is a crucial aspect of conversational e-commerce search. vakulenko2021large revealed the complicated situation for large-scale dialogue analysis specifically focusing on the patterns of mixed-initiative. In contrast with traditional web search scenarios, e-commerce search needs to analyze more complicated user engagement behaviors (See Section 5.1.2 and Section 3.2). Hence in future work, both short-term and long-term user engagement interactions may need specialized attention in conversational e-commerce search. Third, how to evaluate the performance of conversational e-commerce search is still an open question. We have demonstrated a wide-range class of evaluation metrics applied in e-commerce search, e.g., engagement-based metrics and revenue-based metrics (See Section 5.2). Therefore, it is even more challenging to measure the success of conversational e-commerce search in terms of all these various kinds of metrics. Measuring interactivity is critical in conversational information seeking [zamani2022conversational]. Keeping track of how well a user understands the system and vice versa can be another important research direction. Fourth, personalization conversational e-commerce search needs more attention in future work.
@@ 5.5.3 Generative retrieval in e-commerce
Generative retrieval (GR) has emerged as a novel retrieval paradigm in recent years. During the training phase, documents and their corresponding document identifiers (DocIDs) are encoded into the model’s parameters. Given a query, the model can directly generate the relevant DocIDs [tay2022transformer]. This approach leverages the model’s ability to memorize and utilize document representations, enhancing retrieval efficiency and accuracy by bypassing traditional retrieval pipelines.
@@
For indexing strategies, the approach typically involves using the sequence-to-sequence (seq2seq) model to learn the mapping from queries to DocIDs. In addition, a range of tasks have been introduced to enhance the indexing performance, including learning the mapping from documents to DocIDs [tay2022transformer], constructing pseudo-queries for documents and mapping them to DocIDs [wang2022neural], as well as training the model to rank different documents corresponding to the same query [li2024learning].
@@
Document representations are another key component of the GR model. Since DocIDs are the direct output of model, the quality of these representations directly determines the prediction accuracy of the GR model. Existing DocIDs can generally be categorized into two types: numeric-based and text-based.  tay2022transformer introduced three types of numeric DocIDs: (1) unstructured atomic DocIDs, where each document is assigned a random and unique integer identifier, without any structural or semantic information; (2) naively structured string DocIDs, where each document is assigned a random and unique numeric string; and (3) semantically structured DocIDs, which use a hierarchical k-means method, allowing relevant documents to share the same prefix, thereby introducing semantic structure.  wang2022neural proposed that the meaning of a DodID depends on both the position and the prefix context. Therefore, they propose a prefix-aware weight-adaptive decoder to adapt to different DocIDs. sun2024learning proposed learnable document representations by using a discrete autoencoder to encode documents into short, discrete DocIDs. It optimizes these DocIDs by converting documents into DocIDs through an encoder, then training the model to minimize the reconstruction loss of converting DocIDs back to the original documents. Text-based DocIDs are another popular approach, as they can inherit the language model’s text generation ability and do not require learning a new DocIDs vocabulary. The title can be regarded as an intuitive abstract text that represents the content of a document. de2020autoregressive proposed using the title as the DocID and achieved good retrieval performance. However, an arbitrary document may not have an informative and structured title like those in Wikipedia, zhou2022ultron combined keywords from both the URL and the title to form the DocID. Additionally, some works have attempted to incorporate more content into text-based DocIDs, including titles, URLs, document substrings, pseudo-queries, and more [li2023multiview]. The substrings in the document can also represent the information stored in the document. bevilacqua2022autoregressive used arbitrary n-grams in documents as DocID, and retrieves documents under the constrain of a pre-built FM-indexer.
@@
Recommendation methods refer to information filtering techniques, which can be traced back to the 1980s [malone-1987-intelligent], i.e., even before the web was developed. Significant research efforts have been devoted to recommendation in various domains, such as email, movies, books, and music. E-commerce has its unique properties; recommendation methods developed for other domains may not be suitable for e-commerce scenarios.
In this chapter, we focus on recommendation techniques for e-commerce. First, we summarize the key characteristics of e-commerce recommendation, for which a two-stage framework is developed. This forms the mainstream solution for e-commerce recommender systems. We then review models developed for the two stages. We discuss evaluation methodologies of for e-commerce recommendation and conclude with a description of future research directions to help build better e-commerce recommendation solutions.
@@ 6.1 Characteristics of e-commerce recommendation
Large volume of products
The first key characteristic of e-commerce is the large volume of products, which brings challenges to algorithm scalability, as recommendation algorithms need to quickly scan products and select the ones that are of interest to a user.
It is common that an e-commerce platform contains millions of products [kersbergen-2021-learnings]. This sheer number of items inevitably poses a computational challenge to recommender systems. The training of models needs to be efficient to be able to quickly (e.g., hourly or at least daily) refresh the model given new user behavior data and latency needs to be sufficiently low in order to be able expose the large catalog to a large number of users.
Sparsity
The second characteristic is the sparsity of behavior, since a user can only consume (e.g., purchase or click) a few products.
The main aim of e-commerce recommendation is to satisfy the information needs of users in viewing or purchasing products and services. Behavioral data (e.g., browsing, purchases, and clicks) is an important and useful data source for learning the personalized tastes of users. Given the huge volume of products, it is impossible for a user to interact with most items. The products with which a user actually interacts are typically just a small fraction of the entire catalog [li-2021-next-arxiv]. This results in a significant sparsity issue that forces recommendation methods to learn from user behavior on a limited set of products and to generalize their predictions to all other products.
Data richness
The third characteristic is the richness of product and user data (see Chapter 3).
In addition to user behavior data that directly reflects user preferences, rich information about products (e.g., product name, description, categories, images, etc.) and users (e.g., age, gender, occupation, income level, etc.) is available in e-commerce scenarios. Moreover, there is much contextual information associated with user behavior, such as time, location, last purchase, and submitted query in the session etc. [chen-top-n-2017]. This auxiliary information is useful for inferring why a user chooses an item, and is particularly beneficial for cold-start scenarios, where a user (or an item) has very few interactions.
@@
Figure 6.1: Illustration of a two-stage recommendation solution that contains two phases: candidate retrieval and candidate ranking. Image source: [ren2018information].
We present a two-stage solution that has been commonly used in industrial recommender systems [cheng2016wide, Wang2018Improving, Covington:2016]. Fig. 6.1 illustrates a two-stage recommendation framework with two main phases: candidate retrieval and candidate ranking. Given an information need (either expressed explicitly, e.g., by a user expressing interest in a product category, or implicitly, e.g., through the event of a user visiting a recommender system), the first phase of candidate retrieval goes through the entire product catalog, which may contain millions of products, and selects a small set of products that best match the information need. Then the selected candidates (usually in the magnitude of hundreds) are passed to the second phase of candidate ranking, which ranks the candidates to produce the final top-$k$ products to the user. Such a two-stage architecture can strike a balance between efficiency and effectiveness — it not only supports fast retrieval from a large-scale catalog by using an efficient and light-weight candidate retrieval model, but also maintains good recommendation performance as the final ranking is determined using a potentially fine-grained ranking model. We survey studies into each stage in Section 6.2 and Section 6.3, respectively.
It is worth mentioning that the pipeline sketched in Fig. 6.1 may involve third stage, re-ranking, to refine the list produced by the ranking model to meet additional criteria or constraints, such as diversity, novelty, or fairness [abdollahpouri2019popularity, Wilhelm2018Practical, Gogna2017Balancing, Pei2019Personalized, ai2018learning]. We briefly introduce recent progress in re-ranking models in Section 6.4.
@@ 6.2.1 Heuristic methods
Heuristic-based methods are commonly used for candidate retrieval because they are simple and easy to implement. These methods are based on a manually heuristic rather than on optimization with an objective function. Early studies have presented various retrieval strategies to search candidate products. For example, selecting high sale or promotion items has been widely adopted in practical recommender systems, whereas several researchers leveraged principles from economics to perform item selection for candidate retrieval [zhao2017multi, zhang2016economic].
@@
Neighborhood-based methods. Neighborhood-based methods first compute similarity between items or users, and then make recommendations by aggregating information from similar users or items. Neighborhood-based methods can be classified into two types: user-oriented and item-oriented. The paradigm of user-oriented methods [grouplens1994, grouplens1997] can be summarized as follows [adomavicius2005toward]:
@@
where $\hat{r}_{u,i}$ is the rating of user $u$ for the target unrated item $i$ that we seek to estimate, $S_{u}$ is the set of similar users of $u$, and $r_{u^{\prime},i}$ is the rating of similar user $u^{\prime}$ on the target item $i$. $Agg(\cdot)$ is a function that aggregates information from similar users. The process consists of two steps: (i) finding similar users, and (ii) aggregating the information of similar users. Recent work has explored various approaches for similarity computation, including but not limited to Pearson correlation coefficients, information entropy, and mean squared difference [shardanand1995social] between the ratings given by two users [grouplens1994, shardanand1995social]. Spearman rank correlation has been leveraged to measure item rank similarity rather than value similarity [herlocker1999algorithmic]. For the aggregation function $Agg(\cdot)$, a linear weighted combination is a commonly used strategy [grouplens1994, shardanand1995social, herlocker1999algorithmic]. The similarity score is usually leveraged for setting combination weights, as similar users naturally deserve a larger contribution to a prediction.
Inspired by the success of user-oriented methods, item-oriented methods have also been explored [sarwar2001itemCF, karypis2001itemCF]. Item-oriented methods have a quite similar process compared as user-oriented methods except that they aim at mining item similarity rather than user similarity. Recent studies empirically show that item-oriented methods achieve better performance than user-oriented methods [mclaughlin2004useranditemcf, sarwar2001itemCF, karypis2001itemCF].
@@
Complementary and substitutable items. Beyond item similarity, more complicated item relations concerning complementarity of items and substitutability of one item for another have also been considered [Wang2018Improving, zhang2018quality, chen2020try, guoziyi2018]. Complements indicate items that might be purchased together, while substitutes indicate items that are interchangeable. Mining complements and substitutes is beneficial to satisfy the true need of users, and further increases the click-through rate and user stickiness [Wang2018Improving, guoziyi2018]. However, it is challenging as we often lack ground-truth labels of such relations. To deal with this problem, Wang2018Improving leverage co-view and co-purchase statistics, as weak relation signals to supervise the learning of the complements and substitutes. They further integrate the learned relations into a vanilla recommendation model and observe improvements in recommendation performance. guoziyi2018 propose a graph convolutional neural network that decouples item semantics for inferring complementary and substitutable items. The decoupled graph neural network contains a two-step knowledge integration scheme. chen2020try propose attribute-aware collaborative filtering to perform substitute recommendation by addressing issues from both personalization and interpretability perspectives.
Figure 6.3: Illustration of the workflow of embedding-based methods: They first map users and items into a common embedding space; then they leverage the KNN algorithm to search the candidate items having the smallest embedding distance with the target user. Image source: [andoni2006near].

### v3

@@ 5.5.2 Conversational e-commerce search
Originating from early studies on interactive information retrieval [croft1987i3r, belkin1995cases], conversational search refers to the process of interacting with a dialogue system to search for information. Conversational search allows users to express their information needs by directly conducting conversations with search engines. Unlike traditional query-based search engines, conversational search systems capture users’ intent by taking advantage of the flexibility of mixed-initiative interactions and by providing useful information more directly using human-like responses [radlinski-2017-theoretical, Vtyurina:2017:ECS:3027063.3053175, ren2021wizard, vakulenko2021large]. User studies have been conducted to study whether conversational search is needed and what it should look like. Trippas:2018:IDS:3176349.3176387 conduct a laboratory-based observational study, where pairs of people perform search tasks communicating verbally. The authors find that conversation search paradigms are more complex and interactive than traditional search scenarios. Moreover, it is difficult to simulate human-human interactions in a conversational search session. To address this problem, ren2021wizard collect a dataset of human-human dialogues about conversational search in a wizard-of-oz fashion, namely wizard of search engine (WISE), where two workers play the role of seeker and intermediary, respectively.
@@
In a conversational search session, the user first initializes the conversation with a request, then the conversational agent iteratively asks the user about their preferences and estimates user interest based on their feedback. Finally, the agent retrieves the information and generates the response. Many studies focus on the second stage, including generating clarifying queries [Hamedwww2020, liu2021learning, ghanem2022question] and understanding users’ intent [wu-yan-2018-deep, gao-2021-advances, TRIPPAS2020102162]. radlinski-2017-theoretical consider the question of what properties would be desirable for a system so that the system enables users to answer a variety of information needs in a natural and efficient manner. azzopardi2018conceptualizing outline the actions and intents of users and systems and explain how these actions enable users to explore the search space and resolve their information needs. ghanem2022question propose a method to generate user queries for story-based reading comprehension skills. The response generation problem in conversational search is also being addressed. ren2021wizard describe a modular end-to-end neural architecture to transfer the output from intent understanding to improve response generation. ye2022structured design an interconnected network to co-generate structured and natural responses that allow for bidirectional semantic associations to generate responses.
@@
In e-commerce, conversational search systems can help consumers access products through instant conversational interactions on mobile phones or other smart devices. Also, conversational e-commerce search alleviates the burden of reformulating queries and browsing through dozens of products in e-commerce search [zou2022learning]. Moreover, conversational e-commerce search provides a natural way to collect explicit feedback from users to understand their preferences [zamani2022conversational]. As an early attempt, zhang2018towards detail a paradigm for conversational product search to ask users about their preferred values of an aspect and adopt a memory network to retrieve search results. Based on that, bi2019conversational focus on product-seeking conversations and proposed an aspect-value likelihood model for negative feedback, with a multivariate Bernoulli distribution to generate explainable e-commerce aspects. To improve the quality of representation in conversational product search, zou2022learning integrate the representation learning of user, query, item, and conversation into a unified generative framework.
@@
There are several open questions and research directions for future work in conversational e-commerce search. First, obtaining data in real-world scenarios is a challenging problem. Several studies, including, e.g., ConvPS [zou2022learning], have applied simulated data in their experiments. It would be useful to extend observational experiments to a wizard-of-oz setting by establishing a real-world dataset in the future. We also observe that the ongoing trend in simulating users in task-oriented dialogues could provide more useful insights [sunzhang2021simulator]. Second, modeling user-system interactions is a crucial aspect of conversational e-commerce search. vakulenko2021large reveal the complicated situation for large-scale dialogue analysis specifically focusing on the patterns of mixed-initiative. In contrast with traditional web search scenarios, e-commerce search needs to analyze more complicated user engagement behavior (see Section 5.1.2 and 3.2). Hence, in future work, both short-term and long-term user engagement interactions may need specialized attention in conversational e-commerce search. Third, how to evaluate the performance of conversational e-commerce search is still an open question. We have demonstrated a wide range of evaluation metrics applied in e-commerce search, e.g., engagement-based metrics and revenue-based metrics (see Section 5.2). Therefore, it is even more challenging to measure the success of conversational e-commerce search in terms of all these various kinds of metrics. Measuring interactivity is critical in conversational information seeking [zamani2022conversational]. Keeping track of how well a user understands the system and vice versa can be another important research direction. Fourth, personalization conversational e-commerce search needs more attention in future work.
@@ 5.5.3 Generative retrieval in e-commerce
Generative retrieval has emerged as a novel retrieval paradigm in recent years. During the training phase, documents and their corresponding document identifiers (docids) are encoded into the model’s parameters. Given a query, the model can directly generate the relevant DocIDs [tay2022transformer]. This approach uses the model’s ability to memorize and use document representations, enhancing retrieval efficiency and accuracy by bypassing traditional retrieval pipelines.
@@
For indexing strategies, the approach typically involves using a sequence-to-sequence (seq2seq) model to learn the mapping from queries to docids. In addition, a range of tasks have been introduced to enhance the indexing performance, including learning a mapping from documents to docids [tay2022transformer], constructing pseudo-queries for documents and mapping them to docids [wang2022neural], as well as training the model to rank different documents corresponding to the same query [li2024learning].
@@
Document representations are another key component of the generative retrieval model. Since docids are the direct output of model, the quality of these representations directly determines the prediction accuracy of generative models. Existing docids can generally be categorized into two types: numeric-based and text-based. tay2022transformer introduce three types of numeric docids: (i) unstructured atomic docids, where each document is assigned a random and unique integer identifier, without any structural or semantic information; (ii) naively structured string docids, where each document is assigned a random and unique numeric string; and (iii) semantically structured docids, which use a hierarchical $k$-means method, allowing relevant documents to share the same prefix, thereby introducing semantic structure. wang2022neural propose that the meaning of a docid depends on both the position and the prefix context. Therefore, they propose a prefix-aware weight-adaptive decoder to adapt to different docids. sun2024learning detail learnable document representations by using a discrete autoencoder to encode documents into short, discrete docids. It optimizes these docids by converting documents into docids through an encoder, then training the model to minimize the reconstruction loss of converting docids back to the original documents. Text-based docids are another popular approach, as they can inherit the language model’s text generation ability and do not require learning a new docid vocabulary. The title can be regarded as an intuitive abstract text that represents the content of a document. de2020autoregressive describe how to use the document title as a docid and achieve good retrieval performance. However, an arbitrary document may not have an informative and structured title like those in Wikipedia; zhou2022ultron combine keywords from both the URL and the title to form a docid. Additionally, some work has attempted to incorporate more content into text-based docids, including titles, URLs, document substrings, pseudo-queries, and more [li2023multiview]. The substrings in the document can also represent the information stored in the document. bevilacqua2022autoregressive use arbitrary n-grams in documents as docids, and retrieve documents using a pre-built FM-indexer.
@@
Recommendation methods refer to information filtering techniques, which can be traced back to the 1980s [malone-1987-intelligent], i.e., even before the web had been developed. Significant research efforts have been devoted to recommendation in various domains, such as email, movies, books, and music. E-commerce has its unique properties; recommendation methods developed for other domains may not be suitable for e-commerce scenarios.
In this section, we focus on recommendation techniques for e-commerce. First, we summarize the key characteristics of e-commerce recommendation, for which a two-stage framework is developed. This forms the mainstream solution for e-commerce recommender systems. We then review models developed for the two stages. We discuss evaluation methodologies for e-commerce recommendation and conclude with a description of future research directions to help build better e-commerce recommendation solutions.
@@ 6.1 Characteristics of e-commerce recommendation
  • Large volume of products. The first key characteristic of e-commerce is the large volume of products, which brings challenges to algorithm scalability, as recommendation algorithms need to quickly scan products and select the ones that are of interest to a user. It is common that an e-commerce platform contains millions of products [kersbergen-2021-learnings]. This sheer number of items inevitably poses a computational challenge to recommender systems. The training of models needs to be efficient to be able to quickly (e.g., hourly or at least daily) refresh the model given new user behavior data and latency needs to be sufficiently low in order to be able expose the large catalog to a large number of users.
  • Sparsity. The second characteristic is the sparsity of behavior, since a user can only consume (e.g., purchase or click) a few products. The main aim of e-commerce recommendation is to satisfy the information needs of users in viewing or purchasing products and services. Behavioral data (e.g., browsing, purchases, and clicks) is an important and useful data source for learning the personalized tastes of users. Given the huge volume of products, it is impossible for a user to interact with most items. The products with which a user actually interacts are typically just a small fraction of the entire catalog [li-2023-next]. This results in a significant sparsity issue that forces recommendation methods to learn from user behavior on a limited set of products and to generalize their predictions to all other products.
  • Data richness. The third characteristic is the richness of product and user data (see Section 3). In addition to user behavior data that directly reflects user preferences, rich information about products (e.g., product name, description, categories, images, etc.) and users (e.g., age, gender, occupation, income level, etc.) is available in e-commerce scenarios. Moreover, there is much contextual information associated with user behavior, such as time, location, last purchase, and submitted query in the session etc. [chen-top-n-2017]. This auxiliary information is useful for inferring why a user chooses an item, and is particularly beneficial for cold-start scenarios, where a user (or an item) has very few interactions.
@@
Figure 6.1: Illustration of a two-stage recommendation solution with candidate retrieval and candidate ranking. Image source: [ren2018information].
We present a two-stage solution that has been commonly used in industrial recommender systems [cheng2016wide, Wang2018Improving, Covington:2016]. Fig. 6.1 illustrates a two-stage recommendation framework with two main phases: candidate retrieval and candidate ranking. Given an information need (either expressed explicitly, e.g., by a user expressing interest in a product category, or implicitly, e.g., through the event of a user visiting a recommender system), the first phase of candidate retrieval goes through the entire product catalog, which may contain millions of products, and selects a small set of products that best match the information need. Then the selected candidates (usually in the order of hundreds) are passed to the second phase of candidate ranking, which ranks the candidates to produce the final top-$k$ products to the user. Such a two-stage architecture can strike a balance between efficiency and effectiveness – it not only supports fast retrieval from a large-scale catalog by using an efficient and light-weight candidate retrieval model, but also maintains good recommendation performance as the final ranking is determined using a potentially fine-grained ranking model. We survey studies into each stage in Section 6.2 and 6.3, respectively.
It is worth mentioning that the pipeline sketched in Fig. 6.1 may involve a third stage, re-ranking, to refine the list produced by the ranking model to meet additional criteria or constraints, such as diversity, novelty, or fairness [abdollahpouri2019popularity, Wilhelm2018Practical, Gogna2017Balancing, Pei2019Personalized, ai2018learning]. We briefly introduce recent progress in re-ranking models in Section 6.4.
@@ 6.2.1 Heuristic methods
Heuristic-based methods are commonly used for candidate retrieval because they are simple and easy to implement. These methods are based on a manually heuristic rather than on optimization with an objective function. Early studies have presented various retrieval strategies to search candidate products. For example, selecting high sale or promotion items has been widely adopted in practical recommender systems, whereas several researchers use principles from economics to perform item selection for candidate retrieval [zhao2017multi, zhang2016economic].
@@
Neighborhood-based methods. Neighborhood-based methods first compute similarity between items or users, and then make recommendations by aggregating information from similar users or items. Neighborhood-based methods can be classified into user-oriented and item-oriented. The paradigm of user-oriented methods [grouplens1994, grouplens1997] can be summarized as follows [adomavicius2005toward]:
@@
where $\hat{r}_{u,i}$ is the rating of user $u$ for the target unrated item $i$ that we seek to estimate, $S_{u}$ is the set of similar users of $u$, and $r_{u^{\prime},i}$ is the rating of similar user $u^{\prime}$ on the target item $i$. $Agg(\cdot)$ is a function that aggregates information from similar users. The process consists of two steps: (i) finding similar users, and (ii) aggregating the information of similar users. Recent work has explored various approaches for similarity computation, including but not limited to Pearson correlation coefficients, information entropy, and mean squared difference [shardanand1995social] between the ratings given by two users [grouplens1994, shardanand1995social]. Spearman rank correlation has been used to measure item rank similarity rather than value similarity [herlocker1999algorithmic]. For the aggregation function $Agg(\cdot)$, a linear weighted combination is a commonly used strategy [grouplens1994, shardanand1995social, herlocker1999algorithmic]. The similarity score is usually used for setting combination weights, as similar users naturally deserve a larger contribution to a prediction.
Inspired by the success of user-oriented methods, item-oriented methods have also been explored [sarwar2001itemCF, karypis2001itemCF]. Item-oriented methods have a quite similar process as user-oriented methods except that they aim at mining item similarity rather than user similarity. Recent studies empirically show that item-oriented methods achieve better performance than user-oriented methods [mclaughlin2004useranditemcf, sarwar2001itemCF, karypis2001itemCF].
@@
Complementary and substitutable items. Beyond item similarity, more complicated item relations concerning complementarity of items and substitutability of one item for another have also been considered [Wang2018Improving, zhang2018quality, chen2020try, guoziyi2018]. As explained in Section 3.1.2.2, complements indicate items that might be purchased together, while substitutes indicate items that are interchangeable. Mining complements and substitutes is beneficial to satisfy the true need of users, and further increases the click-through rate and user stickiness [Wang2018Improving, guoziyi2018]. However, it is challenging as we often lack ground-truth labels of such relations. To deal with this problem, Wang2018Improving use co-view and co-purchase statistics, as weak relation signals to supervise the learning of the complements and substitutes. They further integrate the learned relations into a vanilla recommendation model and observe improvements in recommendation performance. guoziyi2018 introduce a graph convolutional neural network that decouples item semantics for inferring complementary and substitutable items. The decoupled graph neural network contains a two-step knowledge integration scheme. chen2020try design attribute-aware collaborative filtering to perform substitute recommendation by addressing issues from both personalization and interpretability perspectives.
Figure 6.3: Illustration of the workflow of embedding-based methods: They first map users and items into a common embedding space; then they use the KNN algorithm to search the candidate items having the smallest embedding distance with the target user. Image source: [andoni2006near].